OpenStudy (anonymous):

Barry is working as a pizza delivery boy in south California . He delivers 620 pizzas everyday except on Monday and Tuesdays. If Barry earns 4.2 cents for every second pizza he delivers, how much money can he earn in a month with 28 days?” 378.2$,376.5$,287.8$,260.4$

OpenStudy (anonymous):

Didn't I just answer it?

OpenStudy (ash2326):

in a week he works five days in 4 weeks 20 days money he earns in one day =620/2*4.2=1302 cents or $13.02 per day in a month he works 20 days, his earnings= 20*13.02= $260.4